OpenStudy (anonymous):

If a ping pong ball falls from the top of a building that is 168 feet high, how many seconds will it take for the ball to hit the ground? Remember, the formula for distance is d=16t^2 A. 1 B. 3.2 C. 10.5 D. 16

OpenStudy (anonymous):

plug 168 in for d and solve

OpenStudy (anonymous):

wat

OpenStudy (jziggy):

168 = 16t^2 10.5 = t^2 \[\sqrt(10.5) = t\] Which rounds to approximately 3.2

OpenStudy (anonymous):

yup^^^